OpenStudy (anonymous):

Solve the differential equation\[yu_x+u_y=u\]\[u(x,0)=f(x)\] I got:\[u(x,y)=f(x-\frac{y^2}{2})e^{\frac{y}{2}}\]Anyone care to check this is right?

OpenStudy (anonymous):

how u solve it?

OpenStudy (anonymous):

I believe the problem becomes easier if you divide the original equation by y, so we get:\[u_x+\frac{1}{y}u_y=\frac{1}{y}u\] You assume that there is a function v(s,t) such that:\[v(s,t)=u(x(s,t),y(s,t))\] Then notice that:\[v_t=u_xx_t+u_yy_t\] Choose x_t and y_t so that they equal the coefficients in the original problem so;\[x_t=1\]\[y_t=\frac{1}{y}\]\[v_t=\frac{1}{y}v\] Also since\[u(x,0)=f(x)\]We must have:\[x(s,0)=s\]\[y(s,0)=0\]\[v(s,0)=f(s)\] We now have 3 ordinary differential equations and 3 conditions, solving them: \[x_t=1\]\[x(s,t)=t+h(s)\]\[x(s,0)=s=h(s)\]\[x(s,t)=t+s\] \[y_t=\frac{1}{y}\]\[y(s,t)=\sqrt{2t+2h(s)}\]\[y(s,0)=0=\sqrt{2h(s)}\]\[y(s,t)=\sqrt{2t}\] \[v_t=\frac{1}{y}v\]\[v(s,t)=e^{\frac{t}{y}}h(s)\]\[v(s,0)=f(s)=h(s)\]\[v(s,t)=f(s)e^{\frac{t}{y}}\] Plugging these 3 results back into\[v(s,t)=u(x(s,t),y(s,t))\]Gives:\[f(s)e^{\frac{t}{y}}=u(t+s,\sqrt{2t})\] Using the three results from earlier, lets solve them for s and t. So that:\[t=\frac{y^2}{2}\]\[s=x-\frac{y^2}{2}\] Now we can resubstitute s and t into the equation to put it back in terms of x and y.\[f(x-\frac{y^2}{2})e^{\frac{y}{2}}=u(x,y)\] I hope did that right, I was slightly confused in certain parts. This method is called the method of characterstics, fyi.

OpenStudy (anonymous):

*bookmark i must learn this method

OpenStudy (anonymous):

Yea, unfortunately I couldn't much good info about this method on the internet...

OpenStudy (anonymous):

@eliassaab